OpenStudy (anonymous):

Write a formula for this sequence: 1/5, 3/5, 1, 1 2/5... A. -1/5 + 2/5 n B. 1/5 + 2/5 n C. 1 2/5 - 2/5 n

OpenStudy (anonymous):

b. the n represents any number so that it increases the 1/5 by 2/5. so 1/5+(2/5)(0)=1/5, 1/5+(2/5)(1)= 3/5 and so on